Board Company has a foreign subsidiary that began operations at the start of 2017 with assets of 144,000 kites (the local currency unit) and liabilities of 78,000. During this initial year of operation, the subsidiary reported a profit of 38,000 kites. It distributed two dividends, each for 6,200 kites with one dividend declared on March 1 and the other on October 1. Applicable exchange rates for 1 kite follow:

January 1, 2017 (start of business) $0.81
March 1, 2017 0.79
Weighted average rate for 2017 0.78
October 1, 2017 0.77
December 31, 2017 0.76
  1. Assume that the kite is this subsidiary’s functional currency. What translation adjustment would Board report for the year 2017?

  2. Assume that on October 1, 2017, Board entered into a forward exchange contract to hedge the net investment in this subsidiary. On that date, Board agreed to sell 170,000 kites in three months at a forward exchange rate of $0.77/1 kite. Prepare the journal entries required by this forward contract.

  3. Compute the net translation adjustment for Board to report in accumulated other comprehensive income for the year 2017 under this second set of circumstances.
Answer #1

Net asset ending balance = [(144000-78000) *0.81] + [ 38000*0.78] - [ 6200*0.79] - [6200*0.77] = 73428

Net asset ending bal at current rate = [144000-78000 + 38000 - 6200 - 6200 ] *0.76 = 91600 *0.76 = 69616

Translation adjustment = 73428-69616 = 3812[negative balance]

The journal entries are provided as below:

Date Account Titles Debit Credit
1st October 2017 No Journal Entry Required
31st December 2017 Forward Contract [170,000*(.77-.76)] $1700
Translation Adjustment $1700
(To record adjustment in the value of the forward contract)
31st December 2017 Foreign Currency (Kites) (170,000*.76) $129,200
Cash $129,200
(To record purchase of foreign currency)
31st December 2017 Cash $130900
Foreign Currency (Kites) $129,200
Forward Contract $1700
(To record delivery of foreign currency and closing of forward contract)

net translation adjustment for Board to report = -3812 + 1700 = -2112 [negative balance]